OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution Question 2 (Multiple Choice Worth 1 points) [02.03] Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (texaschic101):

-4 | x + 5 | = -16 --- divide both sides by -4 | x + 5 | = 4 x + 5 = 4 x = 4 - 5 x = -1 -(x + 5) = 4 -x - 5 = 4 -x = 4 + 5 -x = 9 x = -9 x = -1 and x = 9 ================= | 2x + 6 | - 4 = 20 --- add 4 to both sides | 2x + 6 | = 24 2x + 6 = 24 2x = 24 - 6 2x = 18 x = 9 -(2x + 6) = 24 -2x - 6 = 24 -2x = 24 + 6 -2x = 30 x = -15 x = 9 and x = -15 =============== any questions ?